# Learning-based Multi-View Stereo: A Survey

### v1

@@ V-B1 Benchmark Performance
Tab. IV summarizes quantitative results of offline MVS methods. Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better. Since RNN based methods have bad time efficiency, we can find that coarse-to-fine methods become main stream of the community because of their impressive performance and high efficiency in both memory and run-time. In addition, iterative methods based on traditional PatchMatch [40, 103] or RAFT [41, 102, 101] also become popular since they achieve comparable performance as state-of-the-art methods with more lightweight structure. For benchmarks, most learning-based methods perform explicitly better than traditional methods on DTU and Tanks and Temples. However, on ETH3D, traditional methods [189, 127] perform better than all existing learning-based methods. We conjecture that the strong viewpoint variations and large textureless regions in human-made environments still make learning-based methods struggle. Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
V-B2 Memory Consumption and Run-time
@@
Low run-time and memory consumption are essential in most industrial applications with limited computational power and storage, e.g., autonomous driving and AR/VR. Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 31, 35] usually achieve high efficiency in both memory and run-time. In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality. These increase run-time and GPU memory. Recently, many researchers try to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods. We compare and summarize the efficiency of existing efficient methods [117, 32, 39, 40, 102, 41, 99] in Fig. 13. All the methods share similar coarse-to-fine structures to improve efficiency. To further reduce run-time and memory, [40, 102, 41, 99] further get rid of costly 3D convolution and use 2D convolution instead. It can be found many methods [40, 102, 41, 99] can achieve near real-time estimation (about 10 Hz) with relatively high resolution images, e.g., $1152\times 864$. In conclusion, it is promising to further improve both the performance and efficiency for practical applications.
@@
Figure 13: GPU memory consumption and run-time of efficient offline methods [117, 32, 39, 40, 102, 41, 99] on DTU dataset [21]. The image resolution and the number of images are set to $1152\times 864$ and 5 respectively. Experiments are done on a NVIDIA 2080Ti GPU.
V-C Unsupervised MVS with Depth Estimation
@@
We summarize the results of unsupervised MVS methods in Tab. III. Based on the conventional photometric consistency loss (Sec. III-A), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results. Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [42], optical flow consistency [44], neural rendering consistency [43], pseudo-label generation with geometric filtering [132, 45], multi-iteration training [132, 45] and featuremetric loss [45]. It is inspiring that KD-MVS [45] already outperforms its backbone, CasMVSNet [32], and many other fully supervised methods. However, since unsupervised methods mainly test on relatively simple scenes [21, 22], their scalability to more complex and large-scale scenes, e.g., ETH3D [23], remains a question. This good generalization capability across various scenes and datasets is crucial for practical applications. Therefore, evaluating unsupervised MVS methods on various scenes is an important task for future research. Most unsupervised methods use simple MVS backbones [15, 32, 75] and mainly focus on the training strategy. Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
V-D Learning-based MVS without Depth Estimation
V-D1 Voxel-based Methods
@@
We summarize the quantitative results on ScanNet [26] in Tab. V. Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26]. We find that depth map-based MVS methods, e.g., SimpleRecon [35], fuse depth maps with traditional TSDF fusion [64] and can achieve comparable performance as these voxel-based methods with lower complexity. This further shows the advantages of depth map-based methods over voxel-based methods.
TABLE V: Quantitative results of voxel-based MVS methods on ScanNet [26].
@@
TABLE VI: Quantitative results of NeRF-based MVS methods on DTU [21].
@@
Methods GeoNeuS [156] Voxurf [193] NeuS2 [157] PermutoSDF [159] Neuralangelo [158] UniSDF [160]
@@
V-D2 NeRF-based Methods
@@
The quantitative results of nerf-based methods on DTU dataset [21] are summarized in Tab. VI. We visualize the qualitative reconstructions of two representative methods, NeuS [50] and VolSDF [49], in Fig. 14. The most outstanding features of NeuS [50] and VolSDF [49] are that they are self-supervised, i.e., without depth supervision, and produce smooth and complete mesh surfaces. However, many recent methods [156, 154, 155] find that incorporating more explicit geometric supervision explicitly improves the reconstruction quality. Compared with depth map-based MVS methods [25, 15], the main drawback of NeRF-based methods is that they are mostly per-scene optimization problems and need lots of time to train the model on each new scene. Recently, some methods [161, 162, 163] try to make the pipeline generalizable so that there is no need to train the model for new scenes. However, they perform worse than depth map-based methods [15]. Another way to improve training time efficiency is to employ hash grids [146], which also improves surface details. However, these methods [157, 158, 159, 160] need lots of GPU memory and space to train and store the large model. In addition, the scalability of NeRF-based methods remains a problem when attempting to apply on large-scale scenes [23].
Figure 14: Reconstruction results (rendered surface normal map and surface mesh) of NeuS [50] and VolSDF [49] on DTU dataset [21]. NeuS and VolSDF apply different approximation of density with SDF.
V-D3 3D Gaussian Splatting-based Methods
@@
Because of its highly efficient rasterization, 3D Gaussian Splatting [170] becomes popular in novel view synthesis and gradually replaces NeRF [28] in the last half year. Recently, some researchers start exploring the potential of 3D Gaussian Splatting in surface reconstruction. However, there are many challenges. For example, the 3D Gaussians do not correspond well to the actual surface of the scene since the 3D Gaussian has a 3D covariance matrix [52]. Moreover, the surface may contain noisy undulations [52]. However, it is worthwhile to further explore in this direction since 3D Gaussian Splatting is much faster to train than those NeRF-based MVS methods. Moreover, instead of directly extracting the surface from 3D space [52], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a potential direction.
V-D4 Large Feed-forward Methods
@@
As a new trend of 3D reconstruction, large feed-forward methods demonstrate their impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images. Employing a large model that is able to get scaled up, these methods manage to learn strong priors of multi-view geometry from large datasets [30]. However, these methods [53, 179, 54] face limitations when it comes to extending the scale of the scene. For example, the 3D representation employed in LRM [53] is a tri-plane representation [180], which has a small resolution and struggles with scaling to larger sizes due to memory constraints. In addition, the tri-plane representation is difficult to model more complex geometry, e.g., with strong self-occlusions. In addition to tri-plane-based 3D representations, new methods [194, 195] have been proposed that directly reconstruct 3D Gaussians in a feed-forward manner. These new methods demonstrate improved capabilities in reconstructing both objects and scenes, but require posed sparse images as inputs. It is worth noting that the generalization ability of these methods is questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching from plane sweep algorithm from traditional methods. Moreover, these methods typically need high-end GPUs since they are usually GPU memory consuming. It is worth noting that, since large feed-forward methods usually do not enforce geometric constraint explicitly, e.g., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement.
V-E Future Research Directions
V-E1 Datasets and Benchmarks
@@
Recently, there are many researchers trying to solve this problem. TartanAir dataset [196] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments. Precise multi-modal sensor data and ground truth labels are provided, including the stereo RGB image, depth image, segmentation, optical flow, camera poses, and LiDAR point cloud. BlendedMVS [24] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes, shown in Tab. IV. ArKitScenes [197] consists of 5,048 RGB-D sequences, which is more than three times the size of the current largest available indoor dataset, ScanNet. As an extension of ScanNet, ScanNet++ [198] is a large-scale dataset that captures high-quality geometry and color of indoor scenes with high-end laser scanner, DSLR camera, and RGB-D streams from an iPhone. Motivated by ImageNet [199] that drives a remarkable trend of learning from large-scale data in 2D visual tasks, Yu et al. [200] propose MVImgNet, a large-scale dataset of multi-view images collected by shooting videos of real-world objects. MVImgNet demonstrates the potential of various 3D visual tasks, including radiance field reconstruction, multi-view stereo, and view-consistent image understanding. Similarly, Objaverse [30] is a large dataset of objects with 800K+ (and growing) 3D models with descriptive captions, tags, and animations. Both MVImgNet [199] and Objaverse [30] can be used to train large feed-forward models, e.g., LRM [53], for reconstruction, since these models need massive data to learn the strong prior. However, both of them mainly focus on object-level scenes. Large datasets for larger-scale scenes, e.g., rooms, are valuable to be explored.
@@
For evaluation, the main benchmarks [26, 21, 22, 23] are all introduced before 2018. Some of these benchmarks, e.g.,  [21, 22], are currently saturated and the learning-based methods are difficult to further improve the performance on them. Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes. For example, LaMAR [201] is a recent large-scale dataset captured using multiple modern AR devices in diverse environments. It contains challenging short-term appearance and structural changes and high quality LiDAR point cloud ground truth. These make LaMAR a potential benchmark for MVS.
V-E2 View Selection
@@
As discussed in Sec. I-B, view selection is crucial for triangulation quality. Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions. However, view selection is often overlooked and not well studied. For example, all offline depth map-based MVS methods [32, 40, 79, 99] follow MVSNet [15] and use the same simple heuristic strategy to compute scores for neighboring views and sort them. Online depth map-based MVS methods [58, 31, 35] and voxel-based methods [48] also adopt heuristic strategies to choose views with enough pose-distance. Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design of existing methods. Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views. If it is possible to choose the best neighboring views for each pixel, reconstruction quality can be further improved and computation reduced. The learned pixel-wise view weight [63, 40, 41] is one solution for this since it weights the source views differently for each pixel based on its visibility across neighboring views.
V-E3 Depth Fusion
@@
As discussed in Sec. I-D, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps. However, TSDF fusion [64, 65] is not robust enough to the outliers in the depth maps and may have memory issues due to the dense volumetric representation. Depth filtering following Galliani et al. [18] introduces lots of hyper-parameters. Researchers may carefully finetune these hyper-parameters for each scene since it may have a great impact on the evaluation metrics. Therefore, it is valuable to improve the depth fusion step so that the reconstruction quality can be further improved. Recently, there have been some learning-based depth map fusion methods [202, 203]. However, they are limited to small scenes due to the dense volume representation and large memory consumption.
V-E4 Features
@@
One not well-studied topic is what kind of feature extractors are suitable for MVS, as mentioned in Sec. II-A. So far, most of the offline depth map-based methods [15, 32, 78] mainly apply simple 2D CNN and FPN [72] structure as a feature extractor. To improve the receptive fields flexibly, Deformation Convolution [204] is used  [78, 77, 40, 79]). The attention mechanism is also applied in feature learning, where works such as  [90, 79, 77] use intra-attention or inter-attention to capture long-range feature dependencies. Furthermore, MVSFormer [82] digs deeper into using patch-wise ViT [181] as a feature extractor in MVS. The insight is that ViT works better to formulate global feature correlations, and FPN can learn detailed ones. In contrast, online depth map-based methods [31, 35] usually adopt efficient pretrained backbones [73, 70, 74]. Recently, large feed-forward methods [53, 54, 175] adopt large pretrained feature encoders, e.g., ViT [181], DINO [83], to learn strong prior from large-amount of data. However, using a powerful feature extraction network usually increases the computation overhead and reduces the efficiency. Finding a good balance between performance and efficiency is an interesting direction, e.g., distilling the large ViT into a relatively small transformer [205].
V-E5 Real-time and Memory Efficient Models

### v2

@@ 6.2.1 Benchmark Performance
Tab. III summarizes quantitative results of offline MVS methods. Compared with those methods that use direct 3D CNN for regularization, the methods that use RNN and coarse-to-fine regularization perform much better. Since RNN based methods have bad time efficiency, we can find that coarse-to-fine methods become main stream of the community because of their impressive performance and high efficiency in both memory and run-time. In addition, iterative methods based on traditional PatchMatch [39, 102] or RAFT [40, 104, 103] also become popular since they achieve comparable performance as state-of-the-art methods with more lightweight structure. For benchmarks, most learning-based methods perform explicitly better than traditional methods on DTU and Tanks and Temples. However, on ETH3D, traditional methods [189, 123] perform better than all existing learning-based methods. We conjecture that the strong viewpoint variations and large textureless regions in human-made environments still make learning-based methods struggle. Therefore, it is valuable to pay more attention to ETH3D to evaluate the robustness of learning-based methods in real-world scenes.
6.2.2 Memory Consumption and Run-time
@@
Low run-time and memory consumption are crucial in most industrial applications with limited computational power and storage, e.g., robotics and AR/VR. Because of the low-resolution input and the simplicity of network structures, online MVS methods [25, 30, 34] usually achieve high efficiency in both memory and run-time. In contrast, offline MVS methods focus on high-resolution images and usually use computationally expensive network modules, e.g., 3D convolutions, to improve reconstruction quality. These explicitly increase run-time and GPU memory. Recently, many researchers tried to improve the efficiency while maintaining the reconstruction accuracy as other offline MVS methods. We compare the efficiency of state-of-the-art offline MVS methods [31, 39, 104, 40, 100, 79, 111, 76, 86] in Fig. 5. All these methods share similar coarse-to-fine structures to improve efficiency. However, we observe that some methods [31, 79, 111, 76] still require lots of computation resource and are not efficient in either memory or run-time. To further reduce run-time and memory, [39, 104, 40, 100] carefully simplify the architecture, e.g., replacing costly 3D convolution with 2D convolution. Moreover, these methods still achieve competitive performance when compared with the top-performing methods. Therefore, it is promising to further improve both the performance and efficiency for practical applications.
@@
Figure 5: GPU memory consumption and run-time of state-of-the-art offline methods [31, 39, 104, 40, 100, 79, 111, 76, 86] on DTU dataset [21]. The image resolution and the number of images are set to $1600\times 1152$ and 5 respectively. For fair comparison, experiments are done on one workstation with a NVIDIA 2080 Ti GPU.
6.3 Unsupervised MVS with Depth Estimation
@@
We summarize the results of unsupervised MVS methods in Tab. II. Based on the conventional photometric consistency loss (Sec. 4.1), unsupervised MVS methods may have difficulties with many challenging situations, e.g., varying lighting conditions or occlusions, and suffer from downgraded reconstruction results. Therefore, many methods propose to employ different strategies to improve the robustness, e.g., segmentation consistency [41], optical flow consistency [43], neural rendering consistency [42], pseudo-label generation with geometric filtering [128, 44], multi-iteration training [128, 44] and featuremetric loss [44]. It is inspiring that KD-MVS [44] already outperforms its backbone, CasMVSNet [31], and many other fully supervised methods. However, since unsupervised methods mainly test on relatively simple scenes [21, 22], their scalability to more complex and large-scale scenes, e.g., ETH3D [23], remains a question. This good generalization capability across various scenes and datasets is crucial for practical applications. Therefore, evaluating unsupervised MVS methods on various scenes is an important task for future research. Most unsupervised methods use simple MVS backbones [15, 31, 75] and mainly focus on the training strategy. Employing state-of-the-art architectures is also a potential direction that can further boost the performance.
6.4 Learning-based MVS without Depth Estimation
6.4.1 Voxel-based Methods
@@
We summarize the quantitative results on ScanNet [26] in Tab. IV. Since the voxel representation increases the computation overhead explicitly when the scale of scene increases, voxel-based methods mainly focus on indoor scenes, e.g., ScanNet [26]. We find that depth map-based MVS methods, e.g., SimpleRecon [34], fuse depth maps with traditional TSDF fusion [65] and can achieve comparable performance as these voxel-based methods with lower complexity. This further shows the advantages of depth map-based methods over voxel-based methods.
TABLE IV: Quantitative results of voxel-based MVS methods on ScanNet [26].
@@
TABLE V: Quantitative results of NeRF-based MVS methods on DTU [21].
@@
Methods Geo-NeuS [50] Voxurf [193] NeuS2 [148] PermutoSDF [150] Neuralangelo [149] UniSDF [151]
@@
6.4.2 NeRF-based Methods
@@
The quantitative results of nerf-based methods on DTU dataset [21] are summarized in Tab. V. The most outstanding features of NeuS [49] and VolSDF [48] are that they are self-supervised, i.e., without depth supervision, and produce smooth and complete mesh surfaces. However, many recent methods [50, 146, 147] find that incorporating more explicit geometric supervision explicitly improves the reconstruction quality. Compared with depth map-based MVS methods [25, 15], the main drawback of NeRF-based methods is that they are mostly per-scene optimization problems and need lots of time to train the model on each new scene. Recently, some methods [152, 153, 154] try to make the pipeline generalizable so that there is no need to train the model for new scenes. However, they perform worse than depth map-based methods [15]. Another way to improve training time efficiency is to employ hash grids [138], which also improves surface details. However, these methods [148, 149, 150, 151] need lots of GPU memory and space to train and store the large model. In addition, the scalability of NeRF-based methods remains a problem when attempting to apply on large-scale scenes [23].
6.4.3 3D Gaussian Splatting-based Methods
@@
Because of its efficient rasterization, 3D Gaussian Splatting [160] becomes popular in novel view synthesis in the last half year. However, there are many challenges to use reconstruct surfaces with 3DGS. For example, the 3D Gaussians do not correspond well to the actual surface of the scene since the 3D Gaussian has a 3D covariance matrix [51]. Moreover, the surface may contain noisy undulations [51]. However, it is worthwhile to further explore in this direction since 3DGS is much faster to train than those NeRF-based MVS methods. Moreover, instead of directly extracting the surface from 3D space [51], rendering accurate depth maps and then fuse them like depth map-based MVS methods is a promising direction [52, 53]. Additionally, generalizable feed-forward 3DGS methods [194, 195] are interesting since they are proven to improve depth accuracy with rendering loss and do not be trained for each scene.
6.4.4 Large Feed-forward Methods
@@
As a new trend of 3D reconstruction, large feed-forward methods demonstrate impressive performance in challenging settings that other MVS methods cannot handle, e.g., single image input and un-posed sparse images. Employing a large model that is able to get scaled up, these methods manage to learn strong priors of multi-view geometry from large datasets [29]. Since these large feed-forward methods [54, 174, 55] usually do not enforce geometric constraint explicitly, i.e., epipolar geometry, the reconstructed geometry is by default placed in a local coordinate system without absolute scales, making it not feasible for applications requiring precise measurement. Also, these methods reconstruct the target under a globally shared representation, while depth-based MVS is typically performed without global data dependency, making conventional MVS still more suitable for large-scale reconstruction and parallel processing. The massive and costly computation of large feed-forward methods also limit the feasibility of applications.
Benefited from the deep prior rooted in the large datasets, large feed-forwards methods are able to handle some of the typical corner cases, e.g., non-Lambertian surfaces. However it also makes their generalization ability questionable since they learn priors of multi-view geometry completely from the training dataset, while the depth map-based methods utilize feature matching via plane sweep algorithm from traditional methods that are more explainable.
6.5 Future Research Directions
6.5.1 Datasets and Benchmarks
@@
Recently, there are many researchers trying to solve this problem. TartanAir dataset [197] is a large-scale synthetic dataset that is collected in photo-realistic simulation environments. BlendedMVS [24] introduces more large-scale scenes and improves the performance of many MVS methods on real-world scenes, shown in Tab. III. ArKitScenes [198] consists of 5,048 RGB-D sequences, which is more than three times the size of the current largest available indoor dataset, ScanNet. As an extension of ScanNet, ScanNet++ [199] is a large-scale dataset that captures high-quality geometry and color of indoor scenes with high-end laser scanner, DSLR camera, and RGB-D streams from an iPhone. Motivated by ImageNet [200] that drives a remarkable trend of learning from large-scale data in 2D visual tasks, Yu et al. [201] propose MVImgNet, a large-scale dataset of multi-view images collected by shooting videos of real-world objects. MVImgNet demonstrates the potential of various 3D visual tasks, including radiance field reconstruction, multi-view stereo, and view-consistent image understanding. Similarly, Objaverse [29] is a large dataset of objects with 800K+ (and growing) 3D models with descriptive captions, tags, and animations. Both MVImgNet [200] and Objaverse [29] can be used to train large feed-forward models, e.g., LRM [54], for reconstruction, since these models need massive data to learn the strong prior. However, both of them mainly focus on object-level scenes. Large datasets for larger-scale scenes, e.g., rooms, are valuable to be explored.
@@
For evaluation, the main benchmarks [26, 21, 22, 23] are all introduced before 2018. Some of these benchmarks are currently saturated and the methods are difficult to further improve the performance on them. Therefore, it is meaningful to introduce new benchmarks to evaluate the robustness and performance of learning-based methods in challenging real-world scenes. For example, LaMAR [202] is a recent large-scale dataset captured using multiple modern AR devices in diverse environments. It contains challenging short-term appearance and structural changes and high quality LiDAR point cloud ground truth. These make LaMAR a potential benchmark for MVS.
6.5.2 View Selection
@@
As discussed in Sec. 2.2, view selection is crucial for triangulation quality. Picking neighboring views that are suitable for triangulation can not only improve the reconstruction accuracy but also reduce useless computation for the bad views, e.g., views with strong occlusions. However, view selection is often overlooked and not well studied. For example, all offline depth map-based MVS methods [31, 39, 79, 100] follow MVSNet [15] and use the same simple heuristic strategy to compute scores for neighboring views and sort them. Online depth map-based MVS methods [60, 30, 34] and voxel-based methods [47] adopt heuristic strategies to choose views with enough pose-distance. Though it is intractable to incorporate non-differentiable view selection into deep learning, it is worth exploring new view selection strategies since it may improve the reconstruction without changing the model design. Current view selection strategies simply pick the same set of neighboring views for all the pixels in the reference image, and it is probable that different reference pixels have different optimal choices of neighboring views. The learned pixel-wise view weight [94, 39, 40] is one solution for this since it weights the source views differently for each pixel based on its visibility across neighboring views.
6.5.3 Depth Fusion
@@
As discussed in Sec. 2.4, both online and offline MVS methods with plane-sweep use traditional TSDF fusion or depth filtering methods to reconstruct mesh or point cloud from the estimated depth maps. However, TSDF fusion [65, 66] is not robust enough to the outliers in the depth maps and may have memory issues due to the dense volumetric representation. Depth filtering following Galliani et al. [18] introduces lots of hyper-parameters. Researchers may carefully finetune these hyper-parameters for each scene since it may have a great impact on the evaluation metrics. Therefore, it is valuable to improve the depth fusion step so that the reconstruction quality can be further improved. Recently, there have been some learning-based depth map fusion methods [203, 204]. However, they are limited to small scenes due to the dense volume representation and large memory consumption.
6.5.4 Features
@@
One not well-studied topic is what kind of feature extractors are suitable for MVS, as mentioned in Sec. 3.1. So far, most of the offline depth map-based methods [15, 31, 78] mainly apply simple 2D CNN and FPN [72] structure as a feature extractor. To improve the receptive fields flexibly, Deformation Convolution [205] is used  [78, 77, 39, 79]). The attention mechanism is also applied in feature learning, where works such as  [91, 79, 77] use intra-attention or inter-attention to capture long-range feature dependencies. Furthermore, MVSFormer [82] digs deeper into using patch-wise ViT [177] as a feature extractor in MVS. The insight is that ViT works better to formulate global feature correlations, and FPN can learn detailed ones. In contrast, online depth map-based methods [30, 34] usually adopt efficient pretrained backbones [73, 70, 74]. Recently, large feed-forward methods [54, 55, 166] adopt large pretrained feature encoders, e.g., ViT [177], DINO [83], to learn strong prior from large-amount of data. However, using a powerful feature extraction network usually increases the computation overhead and reduces the efficiency. Finding a good balance between performance and efficiency is an interesting direction, e.g., distilling the large ViT into a relatively small transformer [206].
6.5.5 Real-time and Memory Efficient Models